# Improving Discrete Diffusion Unmasking Policies Beyond Explicit Reference Policies

### v1

@@ B.2 Algorithms
20:   Make mini-batch $\{B\}_{i=1}^{b}$ such that $B_{1}\cup B_{2}\cup\dots\cup B_{b}=\{1,L\}$ and $B_{i}\cap B_{j}=\emptyset$ for all $i\neq j$
@@
Specifically, the algorithm is divided into two phases: the sampling phase and the gradient update phase. At the sampling phase, we sample various sequences from $p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})$. In this phase, we gather the extracted feature of the MDM transformer, the token distribution of MDM, the mask token index, and record the probability of the trajectory. At the gradient update phase, we can perform mini-batch training where all the gradients of $g_{\phi}(\cdot|{\bm{x}}_{n})$ in $\mathcal{L}_{\rm{UPO}}$ are independent of others. Furthermore, since we have gathered all the information of $\pi_{\theta}$ that is needed to update the gradient in the sampling phase, we can free the memory of MDM $\theta$. The only parameters requiring forward and backward process in the gradient update phase is the unmasking policy model, which is much smaller than MDM (e.g., LLaDA: 8B, corresponding policy model: 134M). Therefore, we can perform memory-efficient training. For more details, please refer to the algorithm table.
@@ Appendix C Detailed Proof
$p_{g_{\phi_{n}}}({\bm{x}}_{0}|{\mathbf{q}})$ $=\frac{1}{Z_{n-1}({\mathbf{q}})}p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})$
$\cdot\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))\cdot r({\mathbf{q}},{\bm{x}}_{0}))-w_{\epsilon}^{-}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))\cdot(1-r({\mathbf{q}},{\bm{x}}_{0}))))$ (32)
@@
$Z_{n-1}({\mathbf{q}})=r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(-\frac{1}{\beta}(w_{\epsilon}^{-}(r_{g_{\phi_{n-1}}}({\mathbf{q}}))$ (33)
@@
$D_{\rm{KL}}$ $(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}}))-D_{\rm{KL}}(p_{\rm{\theta^{*}}}({\bm{x}}_{0}|{\mathbf{q}})||p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))$ (35)
$=\mathbb{E}_{p_{\theta^{*}}}[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})]-\mathbb{E}_{p_{\theta^{*}}}[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})]$ (36)
$=\mathbb{E}_{p_{\theta^{*}}}[\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})]$ (37)
$=\mathbb{E}_{p_{\theta^{*}}}[\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))\cdot r({\mathbf{q}},{\bm{x}}_{0}))-w_{\epsilon}^{-}(r_{g_{\phi^{*}}}({\mathbf{q}}))\cdot(1-r({\mathbf{q}},{\bm{x}}_{0}))]-\log Z_{\phi^{*}}$ (38)
$=\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))-\log Z_{\phi^{*}}\hskip 28.45274pt\because\forall{\bm{x}}_{0}\in\mathcal{X},p_{\theta^{*}}({\bm{x}}|{\mathbf{q}})>0\rightarrow r({\mathbf{q}},{\bm{x}}_{0})=1$ (39)
$=\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))$
$-\log\Biggr(r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(-\frac{1}{\beta}(w_{\epsilon}^{-}(r_{g_{\phi^{*}}}({\mathbf{q}}))\Biggl)$ (40)
$\geq\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))$
$-\log\Biggr(r_{g_{\rm{ref}}}({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))+(1-r_{g_{\rm{ref}}})({\mathbf{q}})\exp(\frac{1}{\beta}(w_{\epsilon}^{+}(r_{g_{\phi^{*}}}({\mathbf{q}}))\Biggl)$ (41)
$\hskip 170.71652pt\because\forall r\in\{0,1\},\;\;w_{e}^{+}(r)\geq 0,\;w_{e}^{-}(r)\geq 0$
@@ C.2 Surrogate loss of output-level reward maximization
$\mathbb{E}_{\tau}\!\Big[r({\bm{x}}_{0})\,\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ $=\mathbb{E}_{{\bm{x}}_{n},a_{n}}\!\Big[\mathbb{E}\!\big[r({\bm{x}}_{0})\mid{\bm{x}}_{n},a_{n}\big]\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ (tower property)
@@
where we define state-wise value function as the expectation of terminal reward starting from the state, i.e., $V_{g_{\phi}}({\bm{x}}_{n}|{\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:{n-1}}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}},{\bm{x}}_{n})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$ and we denote value function at the initial state as $V_{g_{\phi}}({\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$. Accordingly, state-wise action-value function can be defined as $Q_{g_{\phi}}({\bm{x}}_{n},a_{n}|{\mathbf{q}})=\sum_{{\bm{x}}_{n-1}}\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot V_{g_{\phi}}({\bm{x}}_{n-1}|{\mathbf{q}})$. Please note that Eq. (52)-57 is conceptually equivalent to reversing the derivation process of REINFORCE with baseline (Williams, 1992). Since we want to show that $\nabla_{\phi}\mathcal{L}_{\rm{token}}\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}\nabla_{\phi}\mathcal{L}_{\rm{output}}$, we breakdown $\nabla_{\phi}V_{g_{\phi}}({\mathbf{q}})$ using Lemma 2:
@@
For further explanation, we want to recap $p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{n-1}|{\bm{x}}_{n})=\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})$. Please note that $p_{g_{\phi_{\rm{old}}}}>0$ only if ${\bm{x}}_{n}^{a_{n}}=\mathbb{M}$ and ${\bm{x}}_{n-1}^{a_{n}}\neq\mathbb{M}$ since $g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})=0$ for ${\bm{x}}_{n}^{a_{n}}\neq\mathbb{M}$, $\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})=0$ for all ${\bm{x}}_{n-1}$ satisfying $x_{n-1}^{a_{n}}=\mathbb{M}$ by definition. To simplify the derivation process, we consider $\widehat{\mathcal{L}}_{\rm{token,i}}(\phi,{\mathbf{q}})$ defined as follows:
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigl(\prod_{j=1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigr)\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (65)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (66)
@@
$\hskip 28.45274pt\cdot\Bigr(p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{i-1}|{\bm{x}}_{i},{\mathbf{q}})\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (67)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (68)
@@
$\hskip 28.45274pt\Bigr(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i})\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\Biggr)$ (69)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{i},\dots,{\bm{x}}_{L}\\ a_{i+1},\dots,a_{L}\end{subarray}}\Biggl(\Bigr(\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)$ (70)
@@
$\cdot\sum_{\begin{subarray}{c}{\bm{x}}_{i-1},\dots,{\bm{x}}_{0}\\ a_{i},\dots,a_{1}\end{subarray}}\biggl(\Bigr(\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i})\Bigl)\cdot\Bigr(\prod_{j=0}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}})\Bigl)\cdot r({\mathbf{q}},{\bm{x}}_{0})\biggr)\Biggr)$ (71)

### v2

@@ B.2 Algorithms
20:   Make mini-batch $\{B\}_{i=1}^{b}$ such that $B_{1}\cup B_{2}\cup\dots\cup B_{b}=\{1,\dots,L\}$ and $B_{i}\cap B_{j}=\emptyset$ for all $i\neq j$
@@
Specifically, the algorithm is divided into two phases: the sampling phase and the gradient update phase. At the sampling phase, we sample various sequences from $p_{g_{\phi_{\rm{old}}}}(\cdot|{\mathbf{q}})$. In this phase, we gather the extracted feature of the MDM transformer, the token distribution of MDM, the mask token index, and record the probability of the trajectory. At the gradient update phase, we can perform mini-batch training where all the gradients of $g_{\phi}(\cdot|{\bm{x}}_{n})$ in $\mathcal{L}_{\rm{UPO}}$ are independent of others. Furthermore, since we have gathered all the information of $\pi_{\theta}$ that is needed to update the gradient in the sampling phase, we can free the memory of MDM $\theta$. The only parameters requiring forward and backward process in the gradient update phase is the unmasking policy model, which is much smaller than MDM (e.g., LLaDA: 8B, corresponding policy model: 134M). Therefore, we can perform memory-efficient training. For more details, refer to the algorithm table.
@@ Appendix C Detailed Proof
$p_{g_{\phi_{n}}}({\bm{x}}_{0}\mid{\mathbf{q}})$ $=\frac{1}{Z_{n-1}({\mathbf{q}})}\,p_{g_{\rm ref}}({\bm{x}}_{0}\mid{\mathbf{q}})$ (32)
$\quad\times\exp\!\left(\frac{1}{\beta}\Big[w_{\epsilon}^{+}\!\big(r_{g_{\phi_{n-1}}}({\mathbf{q}})\big)\,r({\mathbf{q}},{\bm{x}}_{0})-w_{\epsilon}^{-}\!\big(r_{g_{\phi_{n-1}}}({\mathbf{q}})\big)\,(1-r({\mathbf{q}},{\bm{x}}_{0}))\Big]\right).$ (33)
@@
$Z_{n-1}({\mathbf{q}})=r_{g_{\rm ref}}({\mathbf{q}})\,\exp\!\left(\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi_{n-1}}}({\mathbf{q}})\big)\right)\;+\;\big(1-r_{g_{\rm ref}}({\mathbf{q}})\big)\,\exp\!\left(-\frac{1}{\beta}\,w_{\epsilon}^{-}\!\big(r_{g_{\phi_{n-1}}}({\mathbf{q}})\big)\right).$ (34)
@@
$D_{\rm{KL}}$ $(p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}}))-D_{\rm{KL}}(p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})\,\|\,p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}}))$ (36)
$=\mathbb{E}_{p_{\theta^{*}}}\!\left[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right]-\mathbb{E}_{p_{\theta^{*}}}\!\left[\log p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})\right]$ (37)
$=\mathbb{E}_{p_{\theta^{*}}}\!\left[\log p_{g_{\phi^{*}}}({\bm{x}}_{0}|{\mathbf{q}})-\log p_{g_{\rm{ref}}}({\bm{x}}_{0}|{\mathbf{q}})\right]$ (38)
$=\mathbb{E}_{p_{\theta^{*}}}\!\Bigg[\frac{1}{\beta}\Big(w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\,r({\mathbf{q}},{\bm{x}}_{0})-w_{\epsilon}^{-}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\,\big(1-r({\mathbf{q}},{\bm{x}}_{0})\big)\Big)-\log Z_{\phi^{*}}\Bigg]$ (39)
$=\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)-\log Z_{\phi^{*}}\hskip 28.45274pt\because\forall{\bm{x}}_{0}\in\mathcal{X},\;p_{\theta^{*}}({\bm{x}}_{0}|{\mathbf{q}})>0\Rightarrow r({\mathbf{q}},{\bm{x}}_{0})=1$ (40)
$=\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)$
$\quad-\log\!\Bigg(r_{g_{\rm{ref}}}({\mathbf{q}})\,\exp\!\Big(\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\Big)+\big(1-r_{g_{\rm{ref}}}({\mathbf{q}})\big)\,\exp\!\Big(-\frac{1}{\beta}\,w_{\epsilon}^{-}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\Big)\Bigg)$ (41)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}>}\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)$
$\quad-\log\!\Bigg(r_{g_{\rm{ref}}}({\mathbf{q}})\,\exp\!\Big(\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\Big)+\big(1-r_{g_{\rm{ref}}}({\mathbf{q}})\big)\,\exp\!\Big(\frac{1}{\beta}\,w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\Big)\Bigg)$ (42)
 $\hskip 56.9055pt{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\because\;1\geq r_{g_{\phi^{*}}}({\mathbf{q}})>r_{g_{\rm ref}}({\mathbf{q}})>0,\;\text{s.t.}\;w_{\epsilon}^{+}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)\geq 0,\;w_{\epsilon}^{-}\!\big(r_{g_{\phi^{*}}}({\mathbf{q}})\big)>0}$
@@
Remark. Why does the support assumption $\operatorname{supp}(p_{\theta^{*}}(\cdot\mid q))\subseteq\operatorname{supp}(p_{g_{\mathrm{ref}}}(\cdot\mid q))$ hold?
Even if $g_{\mathrm{ref}}$ is deterministic (e.g., max-confidence), the support assumption naturally holds for masked diffusion models. The reason is that the MDM denoiser $\pi_{\theta}(\cdot\mid x_{n})$ assigns strictly positive probability to every token in the vocabulary at each unmasking step. Therefore, for any terminal sequence $\mathbf{x}_{0}\in\mathcal{V}^{L}$, there always exists at least one valid denoising trajectory under $g_{\mathrm{ref}}$ that reaches $\mathbf{x}_{0}$, implying $p_{g_{\mathrm{ref}}}(\mathbf{x}_{0}\mid q)>0$. Formally, the terminal-output likelihood marginalizes over all trajectories:
$p_{g_{\mathrm{ref}}}(\mathbf{x}_{0}\mid q)=\sum_{\mathbf{x}_{1:L},\,a_{1:L}}\prod_{n=1}^{L}g_{\mathrm{ref}}(a_{n}\mid\mathbf{x}_{n},q)\,\pi_{\theta}(\mathbf{x}_{n-1}\mid\mathbf{x}_{n},a_{n},q).$
Because $\pi_{\theta}$ has full support, the product above is strictly positive for at least one trajectory. To illustrate, take any target sequence $\mathbf{x}_{0}^{\ast}$. Starting from the fully masked state, $g_{\mathrm{ref}}$ selects an index $a_{L}$; since $\pi_{\theta}(x^{a_{L}}=x_{0}^{\ast,a_{L}}\mid\mathbf{x}_{L})>0$, the next state $\mathbf{x}_{L-1}$ satisfying $x_{L-1}^{a_{L}}=x_{0}^{*,a_{L}}$ is reachable. Repeating this inductively shows that $\mathbf{x}_{0}^{\ast}$ is reachable regardless of $g_{\mathrm{ref}}$. Thus the support assumption holds.
@@ C.2 Surrogate loss of output-level reward maximization
$\mathbb{E}_{\tau}\!\Big[r({\bm{x}}_{0})\,\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ (45)
$=\mathbb{E}_{{\bm{x}}_{n},a_{n}}\!\Big[\mathbb{E}\!\big[r({\bm{x}}_{0})\mid{\bm{x}}_{n},a_{n}\big]\;\nabla_{\phi}\log g_{\phi}(a_{n}\mid{\bm{x}}_{n})\Big]$ (tower property)
@@
where we define state-wise value function as the expectation of terminal reward starting from the state, i.e., $V_{g_{\phi}}({\bm{x}}_{n}|{\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:{n-1}}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}},{\bm{x}}_{n})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$ and we denote value function at the initial state as $V_{g_{\phi}}({\mathbf{q}})\coloneq\mathbb{E}_{{\bm{x}}_{0:L}\sim p_{g_{\phi}}(\cdot\mid{\mathbf{q}})}\bigr[r({\mathbf{q}},{\bm{x}}_{0})\bigl]$. Accordingly, state-wise action-value function can be defined as $Q_{g_{\phi}}({\bm{x}}_{n},a_{n}|{\mathbf{q}})=\sum_{{\bm{x}}_{n-1}}\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot V_{g_{\phi}}({\bm{x}}_{n-1}|{\mathbf{q}})$. Note that Eq. (54)-59 is conceptually equivalent to reversing the derivation process of REINFORCE with baseline (Williams, 1992). Since we want to show that $\nabla_{\phi}\mathcal{L}_{\rm{token}}\mathrel{\vbox{\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}\nabla_{\phi}\mathcal{L}_{\rm{output}}$, we breakdown $\nabla_{\phi}V_{g_{\phi}}({\mathbf{q}})$ using Lemma 2:
@@
For further explanation, we want to recap $p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{n-1}|{\bm{x}}_{n})=\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})\cdot g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})$. Note that $p_{g_{\phi_{\rm{old}}}}>0$ only if ${\bm{x}}_{n}^{a_{n}}=\mathbb{M}$ and ${\bm{x}}_{n-1}^{a_{n}}\neq\mathbb{M}$ since $g_{\phi_{\rm{old}}}(a_{n}|{\bm{x}}_{n})=0$ for ${\bm{x}}_{n}^{a_{n}}\neq\mathbb{M}$, $\pi_{\theta}({\bm{x}}_{n-1}|{\bm{x}}_{n},a_{n})=0$ for all ${\bm{x}}_{n-1}$ satisfying $x_{n-1}^{a_{n}}=\mathbb{M}$ by definition. To simplify the derivation process, we consider $\widehat{\mathcal{L}}_{\rm{token,i}}(\phi,{\mathbf{q}})$ defined as follows:
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}{\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggl(}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl(}\prod_{j=1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr)}\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}r({\mathbf{q}},{\bm{x}}_{0}){\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggr)}$ (67)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}{\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggl(}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}$
@@
$\hskip 28.45274pt\cdot{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{i-1}|{\bm{x}}_{i},{\mathbf{q}})\frac{g_{\phi}(a_{i}\mid{\bm{x}}_{i})}{g_{\phi_{\rm old}}(a_{i}\mid{\bm{x}}_{i})}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=1}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot r({\mathbf{q}},{\bm{x}}_{0}){\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggr)}$ (68)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{0},\dots,{\bm{x}}_{L}\\ a_{1},\dots,a_{L}\end{subarray}}{\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggl(}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}$
@@
$\hskip 28.45274pt{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=1}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot r({\mathbf{q}},{\bm{x}}_{0}){\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggr)}$ (69)
@@
$=\sum_{\begin{subarray}{c}{\bm{x}}_{i},\dots,{\bm{x}}_{L}\\ a_{i+1},\dots,a_{L}\end{subarray}}{\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggl(}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=i+1}^{L}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}$
@@
$\cdot\sum_{\begin{subarray}{c}{\bm{x}}_{i-1},\dots,{\bm{x}}_{0}\\ a_{i},\dots,a_{1}\end{subarray}}{\color[rgb]{0,0.62890625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62890625,0}\biggl(}{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\pi_{\theta}({\bm{x}}_{i-1}|{\bm{x}}_{i},a_{i},{\mathbf{q}})g_{\phi}(a_{i}\mid{\bm{x}}_{i}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot{\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigr(}\prod_{j=1}^{i-1}p_{g_{\phi_{\rm{old}}}}({\bm{x}}_{j-1}|{\bm{x}}_{j},{\mathbf{q}}){\color[rgb]{1,0.55078125,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.55078125,0}\bigl)}\cdot r({\mathbf{q}},{\bm{x}}_{0}){\color[rgb]{0,0.62890625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62890625,0}\biggr)}{\color[rgb]{0.1171875,0.3515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.3515625,1}\Biggr)}$ (70)